OpenStudy (irish_waffl3):

if m(x)=5x+4, and n(x)=6x-9, what is m[n(x)] will fan and medal :)

OpenStudy (yoongibear):

@.Mehek.

OpenStudy (irish_waffl3):

hayulp!

OpenStudy (yoongibear):

im not good at math, mate ;-; sorry

OpenStudy (irish_waffl3):

ok, thanks anyway!

OpenStudy (irish_waffl3):

@.Mehek.

OpenStudy (yoongibear):

i tried to get @.Mehek. here bc she's really smart, but oh well. Im sorry I can't help wow my english is on point XD

OpenStudy (irish_waffl3):

ok XD!

OpenStudy (irish_waffl3):

@welshfella

OpenStudy (yoongibear):

wait

OpenStudy (irish_waffl3):

ok

OpenStudy (yoongibear):

@nincompoop

OpenStudy (yoongibear):

@Jeneshisu notice me senpai

OpenStudy (irish_waffl3):

lol!

OpenStudy (jeneshisu):

I have come. *strange moist smoke apears*

OpenStudy (yoongibear):

ew no xD

OpenStudy (yoongibear):

thanks though, i guess. are you good at math?

OpenStudy (jeneshisu):

I'll try

OpenStudy (yoongibear):

i tried and it fried my brain XD

OpenStudy (jeneshisu):

I can't just type the answer right?

OpenStudy (yoongibear):

nope

OpenStudy (yoongibear):

we're suppose to help them understand so they can do it on their own

OpenStudy (yoongibear):

something like that

OpenStudy (jeneshisu):

oh yea yeah

OpenStudy (irish_waffl3):

so can u help?

OpenStudy (jeneshisu):

give me a sec

OpenStudy (irish_waffl3):

kk!

OpenStudy (jeneshisu):

You just have to sub in the x in nx into the mx. That won't make sense right?

OpenStudy (jeneshisu):

Just substitute the nx into the parentheses and then show me. m[n(x)]

OpenStudy (yoongibear):

i don't even understand e.0

OpenStudy (jeneshisu):

This is what it will look like once you substitute everything in. 5(6x-9)+4=

OpenStudy (jeneshisu):

It's really easy i just can't explain

OpenStudy (jeneshisu):

@irish_waffl3 replyy plz

OpenStudy (yoongibear):

@irish_waffl3 u there m8?

OpenStudy (yoongibear):

@.Mehek. ur help is needed, oh wise one

OpenStudy (jeneshisu):

where he at ...? Im over here stressin out on how to explain this and he not even here (┛◉Д◉)┛彡┻━┻

OpenStudy (yoongibear):

:o

OpenStudy (yoongibear):

maybe their brain fried o.0

OpenStudy (yoongibear):

and now they're just staring at the screen

OpenStudy (yoongibear):

that would be me tbh

OpenStudy (jeneshisu):

prolly

OpenStudy (jeneshisu):

lol

OpenStudy (irish_waffl3):

oh im back sorry

OpenStudy (irish_waffl3):

@.Mehek.

OpenStudy (jeneshisu):

thank goodness

OpenStudy (jeneshisu):

Did you get what I wrote above...? ( ´_⊃`)

OpenStudy (irish_waffl3):

yep! 5(6x-9)+4=

OpenStudy (jeneshisu):

Ok now you must do the distributive property and tell me what you get

OpenStudy (irish_waffl3):

ok

OpenStudy (irish_waffl3):

5(6x)-45+4 ??

OpenStudy (jeneshisu):

nope, you have to multiply 5 and 6x, and 5 times -9

OpenStudy (irish_waffl3):

5*-9 is -45

OpenStudy (irish_waffl3):

5*6x is not 30x, so what is it?

OpenStudy (jeneshisu):

Wym it's not?

OpenStudy (irish_waffl3):

it is nvm

OpenStudy (irish_waffl3):

i gots confused

OpenStudy (irish_waffl3):

so the answer is 30x-49?

OpenStudy (jeneshisu):

no hold on

OpenStudy (jeneshisu):

now you have 30x-45+4

OpenStudy (jeneshisu):

that's your answer.

OpenStudy (irish_waffl3):

but wouldn't you add 45+4?

OpenStudy (irish_waffl3):

making it m[n(x)=30x+49?

OpenStudy (irish_waffl3):

@Jeneshisu ?

OpenStudy (jeneshisu):

it would be -45+4 =

OpenStudy (irish_waffl3):

oh yeh

OpenStudy (irish_waffl3):

so -41

OpenStudy (jeneshisu):

Yupp!

OpenStudy (irish_waffl3):

thanx!

OpenStudy (jeneshisu):

anytime

OpenStudy (irish_waffl3):

as i promised, i medaled and fanned